40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9667     Entry time: Mon Feb 24 23:43:10 2014
Author: rana 
Type: Summary 
Category: General 
Subject: ToDo 

1) Fixup REFL165: remove ND filters, get box for PD, dump diode reflections, put less light on diode, change DC transimpedance (?), max power dissipation on BBPD < 0.5 W w/ 25 V bias. Perhaps replace OP27 with TLE2027.

2) Make plan for fixing fiber layout up and down the arms. Need tubing for the whole run. Don't make it cheesy. Two fibers per arm.

3) Fix LSC model to allow user switching of whitening. Get back to working on AutoLock scripts (not Guardian).

4) Manasa, Q, Jenne, tune Oplev servos Tuesday morning/afternoon.

5) Reconnect the other seismometers (Steve, Jenne). For real.

6) Balance PRMI coils at high frequency.

ELOG V3.1.3-